# references.bib

### v3

@@ Abstract
In this work, we aim to reconcile several apparently contradictory observations in market microstructure: is the famous “square-root law” of metaorder impact that decays with time compatible with the random-walk nature of prices and the linear impact of order imbalances? Can one entirely explain the volatility of prices as resulting from the flow of uninformed metaorders that mechanically impact prices? We introduce a new theoretical framework to describe metaorders with different signs, sizes and durations, which all impact prices as a square-root of volume but with a subsequent time decay. We show that, as in the original propagator model, price diffusion is ensured by the long memory of cross-correlations between metaorders. In order to account for the effect of strongly fluctuating volumes $q$ of individual trades, we further introduce two $q$-dependent exponents, which allows us to account for the way the moments of generalized volume imbalance and the correlation between price changes and generalized order flow imbalance scales with $T$. We predict in particular that the corresponding power-laws depend in a non-monotonic fashion on a parameter $a$ that allows one to put the same weight on all child orders or overweight large orders, a behaviour clearly borne out by empirical data. We also predict that the correlation between price changes and volume imbalances should display a maximum as a function of $a$, which again matches observations. Such noteworthy agreement between theory and data suggests that our framework correctly captures the basic mechanism at the heart of price formation, namely the average impact of metaorders. We argue that our results support the “Order-Driven” theory of excess volatility, and are at odds with the idea that a “Fundamental” component accounts for a large share of the volatility of financial markets.
@@ 1 Introduction
Another, very different interpretation of price impact is that it is a purely statistical reaction of the market to incoming order flow, where information plays little role. Prices move just because people trade, whatever the reason they are trading, and volatility is the result of people randomly buying and selling. This is the Order-Driven view of markets, explicitly spelled out in [bouchaud2018trades], chapter 20 – but see also [bouchaud2006random, hopman2007supply, bouchaud2010impact] and in a different setting, [gabaix2021search, bouchaud2022inelastic]. In this scenario, there is no “information revelation” but rather “self-fulfilling prophecies”, as recently emphasized in [van2024ponzi].
Of course, reality should lie somewhere between these two extremes. There are surely some informed trades, and news do impact prices, but there is also overwhelming evidence for the presence of “noise traders”, excess trading and excess volatility in financial markets, see e.g. [black1986noise, odean1999investors, shiller1981stock, leroy2006excess, kurth2025revisiting]. Direct empirical estimates suggest that informed trades are a minority (see [bouchaud2018trades], chapter 16). While we are convinced that the Order-Driven view is a much closer approximation to the dynamics of markets, there are several empirical loose ends that need to be tied up. A major issue is how to reconcile two prominent stylized facts of market microstructure, namely, (i) the long-range memory in order signs (i.e. $+1$ for buy orders and $-1$ for sell orders) and (ii) the ubiquitous square-root law of market impact (that governs the average price move induced by the execution of a sequence of orders) with the random walk nature of prices, with a volatility $\sigma$ that is directly proportional to the amplitude of the square-root law. More precisely, the square-root law states that the average price impact $\mathcal{I}$ of a metaorder of total volume $Q$ is given by
@@
Eq. (1) may look familiar, and in fact trivial: superficially, it states that price changes (i.e. $\mathcal{I}(Q)$) grow as the square-root of execution time $T$, i.e. the law of random walks, provided one assumes that $T$ and $Q$ are proportional. But, as argued in [bucci_impact_vol2019], such an argument is completely misleading: not only $\mathcal{I}(Q)$ is an average price change and not a standard deviation; but also $\mathcal{I}(Q)$ is found to depend only on the quantity executed $Q$ and not on execution time $T$. Furthermore, it is known that the impact of metaorders decays, on average, after the end of the execution period (e.g. [moro2009market, brokmann2015slow, bucci2018slow]). Note that all these features are at odds with the Kyle model, that predicts linear and permanent impact, resulting from information revelation [kyle1985continuous].
@@ 3.4.1 Child volume distribution
As discussed in Section 3.2, we consider a log-normal distribution for the child order sizes, as defined by Eq. (28). It turns out to be a reasonable approximation of reality for large volumes, see Fig. 1, with values of $\sigma_{\ell}$ reported in the legend, around $1$ for stocks and SPMINI and $1.2$ for EUROSTOXX. A better fit of the tail of the distribution is, arguably, power-law $\propto q^{-1-\varpi}$, with $\varpi$ found to be around 2.12.12.12.12.42.42.42.4. Such a choice makes the mathematical analysis more cumbersome, and we prefer the log-normal specification for our semi-quantitative discussion of the role of volume fluctuations. Nevertheless, a power-law tail can play a role similar to the coefficient $\lambda$ in $\mu_{q}$ when it comes to the scaling of $\Sigma^{(2n)}_{a}$, crossing over to a linear behaviour when $2na\gtrsim\varpi$, see footnote 2.
@@ 4 The Impact-Diffusivity puzzle and a generalized propagator
We now discuss the problem of whether the volatility of price changes can be explained chiefly in terms of the impact of intertwined metaorders of different sizes and signs that get executed in the market. We will first recall how price diffusivity and long-memory of order flow are reconciled within the standard “propagator” framework, and then show how the issue becomes much more perplexing when the impact of metaorders obeys the square-root law. Three possible resolutions are proposed, together with their strengths and weaknesses.
@@ 4.1 Price diffusivity within the propagator model
Price variations are then given by the sum of two terms, $\Delta_{T,1}$ describing the impact of metaorders initiated within $[0,T]$, and $\Delta_{T,2}$ the decaying impact of metaorders initiated before $t=0$:
@@ 4.2 A generalized propagator model
which is again sub-diffusive whenever $\beta>0$ and $\gamma=\mu-1>0$. In other words, price diffusion is only possible when impact is permanent, i.e. $\beta=0$. This is in fact the assumption made by Sato & Kanasawa in their latest paper [sato2025exactlysolvablemodelsquareroot]. However, all empirical data known to us suggest that impact decays, at least over short to medium time scales [moro2009market, brokmann2015slow, bucci2018slow, maitrier2025generatingrealisticmetaorderspublic], which according to our calculation should lead to substantial price mean reversion on such time scales.
@@ 4.3 The role of metaorder autocorrelations
$\sigma^{2}=C(\beta,\gamma_{\times})\Gamma\nu^{2}\tau_{0}^{\gamma_{\times}}\mathcal{I}_{1}^{2}(q,\varphi)\left(\overline{s^{\frac{1}{2}+\beta}}\right)^{2},$ (57)
where $C(\beta,\gamma_{\times})$ is a numerical coefficient found to be $\approx 5.6$ for $\gamma_{\times}=0.6$ and $\beta=0.2$
@@
$\theta_{0}=Y\frac{\sigma}{\sqrt{\phi}},\qquad Y\propto\frac{\bar{n}^{\frac{1}{2}-\beta}}{\sqrt{C\Gamma}},$ (58)
@@ 4.4 The role of volume fluctuations
Taking $\theta(q)=\theta_{0}\sqrt{q}$ one gets $\mathcal{I}_{1}(q>q_{0},\varphi)=2\theta_{0}\sqrt{\varphi q}$, so the long-term price volatility is given by
@@
We see that this result suggests a weak dependence of the prefactor of the square-root law in $q$ and $\widetilde{\varphi}$, which disappears for large enough $q>q_{0}$, since $\beta_{q}\to 0$ in that case. However, in that case $Y$ would be substantially larger than empirically observed, since from Fig. 3 we estimate $\phi_{0}\approx 0.1\phi$. Besides, since child orders of size $<q_{0}$ (which are the most numerous) only impact prices temporarily, this scenario would lead to strong visible mean-reversion in prices not observed in data – see e.g. Fig. 6 for $n=1$.
@@ 4.6 Discussion
We thus have three possible scenarios for generating long term volatility from the impact of metaorders. The idea that metaorders are correlated between one another, with roughly the same long memory as within each one of them, seems plausible, is compatible with available data and provides a natural extension of the propagator scenario: diffusive prices emerge from the subtle interplay between decaying impact and autocorrelated order flow. Such a scenario predicts a weak dependence of the prefactor $Y$ of the square-root law with the participation rate of the metaorder, as $\widetilde{\varphi}^{\beta}$ with $\beta\approx 0.2$.
@@
Finally, long term volatility may result from the random component of impact, assumed to be permanent. This bypasses the paradox that average impact decays and, in the absence of correlations between metaorders, should lead to subdiffusion. In such a scenario, volatility is induced by trading activity alone, even if average impact was zero. Still, the fact that impact fluctuations are proportional to average impact, as postulated in Eq. (68), is important to recover the correct relation between peak impact $\mathcal{I}(Q)$ and $\sigma\sqrt{Q/\phi}$, Eq. (73), which is now independent of $\widetilde{\varphi}$.
We now turn to the study of the covariance of price changes and volume imbalances, to see whether we can constrain the theory further. We will indeed see that the data strongly favors an interpretation based on the average impact of correlated metaorders. The last scenario, where volatility arises from the fluctuating part of impact, does not pass the test.
@@ 5 Covariance between order flow imbalance and prices changes
Another quantity that can be computed within our model and easily measured empirically using the public tape of trades and prices is the “aggregated” impact $\mathbb{E}[\Delta|I^{a}]$, conditioned to a certain value of imbalance $I^{a}$, as studied in [plerou2002quantifying, patzelt2018universal]. We know that such a quantity behaves very differently from the square-root law, and has non-trivial scaling properties as a function of $T$, see [patzelt2018universal] for $a=0$ and $a=1$. For $a=0$, in particular, one finds that the initial slope of $\mathbb{E}[\Delta|I^{0}]$ as a function of $I^{0}$ scales like $T^{-\omega}$ with $\omega\approx 1/4$ [patzelt2018universal, bouchaud2018trades], a result we confirm in Fig. 7.
@@ 5.5 With “informed” metaorders
Up to now, we assumed that there are no correlations between the sign of metaorders $\varepsilon$ and the “Fundamental” component of price changes on time scale $T$, $\sigma_{F}\xi\sqrt{T}$, see Eq. (68). If we rather assume that $\mathbb{E}[\varepsilon\xi]=\rho/\sqrt{\nu T}$, where $\rho$ measures the average amount of information of individual metaorders, [^11] For a detailed discussion of the scaling with $T$, see [bouchaud2010impact], section 16.1.3. The idea is that out of $N_{T}=\nu T$ metaorders of random signs, an excess fraction $\sim\sqrt{\nu T}$ is possibly informed. It is important to stress that, in the spirit of the Kyle model [kyle1985continuous], the fundamental component $\sigma_{F}\xi\sqrt{T}$ is not a mechanical consequence of impact. we find an extra contribution to $\mathbb{E}[\Delta_{T}\cdot I_{T}^{a}]$ that reads:
@@ 5.6 The correlation coefficient
For $a\to 0$, the power of $T$ is very close to zero for our default choice of parameters. For higher values of $a$, the exponent becomes positive and therefore this off-diagonal contribution adds an increasing function of $T$ [^12] In the $T\to\infty$ limit, one should take into account the fact that $\Sigma_{a}$ itself becomes dominated by the off-diagonal contribution, see the discussion around Eq. (39). Hence the correlation does saturate when $T\to\infty$, as it should be.
@@ 5.7 Empirical data
    Provided the main source of price moves is the average impact of random metaorders, the covariance $\operatorname{\mathbb{E}}[\Delta_{T}\cdot I_{T}^{a}]$ behaves as a power-law of $T$ with an effective exponent that is non-monotonic in $a$, reaching a minimum for some value of $a$, see Eq. (80). If volatility is dominated by the random component of impact (Eq. (68)) with a “Fundamental” component unrelated to trading, a linear behaviour in $T$ independent of $a$ should be observed;
@@ 5.7.4 Empirical covariance: conclusion
  • Stocks and futures seem to differ quantitatively when it comes to the correlation between order imbalance and price changes. In particular, the correlation between the two is stronger for futures contract, and is reached for longer time intervals $T$ and with more weight given on large child order volumes (i.e. $a^{\star}=1$ instead of $a^{\star}=1/2$). This points towards a stronger role of metaorder correlations for futures than for stocks.
  • The hypothesis according to which most of the volatility comes from fundamental information is hard to reconcile with the data. First, as discussed in section 4.5, this assumption does not enable one to understand why the square-root law, which applies to all metaorders – informed or not [gomes2015market, bucci2018slow, sato2024does] – is proportional to volatility. Second, it predicts that the covariance of price returns and volume imbalances is proportional to $T$ independently of $a$, at odds with empirical data. As argued in [bouchaud2006random, taranto2018linear, bouchaud2018trades], the most plausible hypothesis is that volatility stems from trading alone. In other words, the excess volatility puzzle has a microstructural origin.

### v4

@@ Abstract
In this work, we aim to reconcile several apparently contradictory observations in market microstructure: is the famous “square-root law” of metaorder impact, which decays with time, compatible with the random-walk nature of prices and the linear impact of order imbalances? Can one entirely explain the volatility of prices as resulting from the flow of uninformed metaorders that mechanically impact them? We introduce a new theoretical framework to describe metaorders with different signs, sizes and durations, which all impact prices as a square-root of volume but with a subsequent time decay. We show that, as in the original propagator model, price diffusion is ensured by the long memory of cross-correlations between metaorders. In order to account for the effect of strongly fluctuating volumes $q$ of individual trades, we further introduce two $q$-dependent exponents, which allow us to describe how the moments of generalized volume imbalance and the correlation between price changes and generalized order flow imbalance scale with $T$. We predict in particular that the corresponding power-laws depend in a non-monotonic fashion on a parameter $a$, which allows one to put the same weight on all child orders or to overweight large ones, a behaviour that is clearly borne out by empirical data. We also predict that the correlation between price changes and volume imbalances should display a maximum as a function of $a$, which again matches observations. Such noteworthy agreement between theory and data suggests that our framework correctly captures the basic mechanism at the heart of price formation, namely the average impact of metaorders. We argue that our results support the “Order-Driven” theory of excess volatility, and are at odds with the idea that a “Fundamental” component accounts for a large share of the volatility of financial markets.
@@ 1 Introduction
Another, very different interpretation of price impact is that it is a purely statistical reaction of the market to incoming order flow, where information plays little role. Prices move simply because people trade, whatever the reason they are trading, and volatility is the result of people randomly buying and selling. This is the Order-Driven view of markets, explicitly spelled out in [bouchaud2018trades], chapter 20 – but see also [bouchaud2006random, hopman2007supply, bouchaud2010impact] and in a different setting, [gabaix2021search, bouchaud2022inelastic]. In this scenario, there is no “information revelation” but rather “self-fulfilling prophecies”, as emphasized in [bouchaud2009markets, van2024ponzi].
Of course, reality should lie somewhere between these two extremes. There are surely some informed trades, and news do impact prices, but there is also overwhelming evidence for the presence of “noise traders”, excess trading and excess volatility in financial markets, see e.g. [black1986noise, odean1999investors, shiller1981stock, leroy2006excess, kurth2025revisiting]. Direct empirical estimates suggest that informed trades are a minority (see [bouchaud2003fluctuations] and [bouchaud2018trades], chapter 16). While we are convinced that the Order-Driven view is a much closer approximation to the dynamics of markets, there are several empirical loose ends that need to be tied up. A major issue is how to reconcile two prominent stylized facts of market microstructure, namely, (i) the long-range memory in order signs (i.e. $+1$ for buy orders and $-1$ for sell orders) and (ii) the ubiquitous square-root law of market impact (that governs the average price move induced by the execution of a sequence of orders) with the random walk nature of prices, with a volatility $\sigma$ that is directly proportional to the amplitude of the square-root law. More precisely, the square-root law states that the average price impact $\mathcal{I}$ of a metaorder of total volume $Q$ is given by
@@
Eq. (1) may look familiar, and in fact trivial: superficially, it states that price changes (i.e. $\mathcal{I}(Q)$) grow as the square-root of execution time $T$, i.e. the law of random walks, provided one assumes that $T$ and $Q$ are proportional. But, as argued in [bucci_impact_vol2019], such an argument is completely misleading: not only $\mathcal{I}(Q)$ is an average price change and not a standard deviation; but also $\mathcal{I}(Q)$ is found to depend only on the quantity executed $Q$ and not on execution time $T$. Furthermore, it is known that the impact of metaorders decays, on average, after the end of the execution period (e.g. [moro2009market, brokmann2015slow, bucci2018slow]). Note that all these features are at odds with the Kyle model that predicts linear and permanent impact, resulting from information revelation [kyle1985continuous]. [^1] Related to this discussion, see also [saddier2024bayesian].
@@ 3.4.1 Child volume distribution
As discussed in Section 3.2, we consider a log-normal distribution for the child order sizes, as defined by Eq. (28). It turns out to be a reasonable approximation of reality for large volumes, see Fig. 1, with values of $\sigma_{\ell}$ reported in the legend, around $1$ for stocks and SPMINI and $1.2$ for EUROSTOXX. A better fit of the tail of the distribution is, arguably, power-law $\propto q^{-1-\varpi}$, with $\varpi$ found to be around 2.12.12.12.12.42.42.42.4. Such a choice makes the mathematical analysis more cumbersome, and we prefer the log-normal specification for our semi-quantitative discussion of the role of volume fluctuations. Nevertheless, a power-law tail can play a role similar to the coefficient $\lambda$ in $\mu_{q}$ when it comes to the scaling of $\Sigma^{(2n)}_{a}$, crossing over to a linear behaviour when $2na\gtrsim\varpi$, see footnote 3.
@@ 4 The Impact-Diffusivity puzzle and a generalized propagator
We now discuss the problem of whether the volatility of price changes can be explained chiefly in terms of the impact of intertwined metaorders of different sizes and signs that get executed in the market, i.e. without a “Fundamental” component that makes prices move without any trades. We will first recall how price diffusivity and long-memory of order flow are reconciled within the standard “propagator” framework, and then show how the issue becomes much more perplexing when the impact of metaorders obeys the square-root law. Three possible resolutions are proposed, together with their strengths and weaknesses.
@@ 4.1 Price diffusivity within the propagator model
Neglecting any “Fundamental” contributions, price variations are then given by the sum of two terms, $\Delta_{T,1}$ describing the impact of metaorders initiated within $[0,T]$, and $\Delta_{T,2}$ the decaying impact of metaorders initiated before $t=0$:
@@ 4.2 A generalized propagator model
which is again sub-diffusive whenever $\beta>0$ and $\gamma=\mu-1>0$. In other words, impact-driven price diffusion is only possible when impact is permanent, i.e. $\beta=0$. This is in fact the assumption made by Sato & Kanasawa in their latest paper [sato2025exactlysolvablemodelsquareroot]. However, all empirical data known to us suggest that impact decays, at least over short to medium time scales [moro2009market, brokmann2015slow, bucci2018slow, maitrier2025generatingrealisticmetaorderspublic], which according to our calculation should lead to substantial price mean reversion on such time scales.
@@ 4.3 The role of metaorder autocorrelations
$\sigma^{2}=C(\beta,\gamma_{\times})\Gamma\nu^{2}\tau_{0}^{\gamma_{\times}}\mathcal{I}_{1}^{2}(q,\varphi)\left(\overline{s^{\frac{1}{2}+\beta}}\right)^{2}+\sigma^{2}_{F},$ (57)
where $C(\beta,\gamma_{\times})$ is a numerical coefficient found to be $\approx 5.6$ for $\gamma_{\times}=0.6$ and $\beta=0.2$, and $\sigma_{F}$ is the possible “Fundamental” component of volatility (i.e. news driven).
@@
$\theta_{0}=Y\frac{\sigma}{\sqrt{\phi}},\qquad Y\propto\frac{\bar{n}^{\frac{1}{2}-\beta}}{\sqrt{C\Gamma}}\times\sqrt{1-\frac{\sigma_{F}^{2}}{\sigma^{2}}}$ (58)
@@ 4.4 The role of volume fluctuations
Taking $\theta(q)=\theta_{0}\sqrt{q}$ one gets $\mathcal{I}_{1}(q>q_{0},\varphi)=2\theta_{0}\sqrt{\varphi q}$, so the contribution to long-term price volatility is given by
@@
We see that this result suggests a weak dependence of the prefactor of the square-root law in $q$ and $\widetilde{\varphi}$, which disappears for large enough $q>q_{0}$, since $\beta_{q}\to 0$ in that case. However, in that case $Y$ would be substantially larger than empirically observed, since from Fig. 3 we estimate $\phi_{0}\approx 0.1\phi$. Besides, since child orders of size $<q_{0}$ (which are the most numerous) only impact prices temporarily, this scenario would again lead to strong visible mean-reversion in prices not observed in data – see e.g. Fig. 6 for $n=1$.
@@ 4.6 Discussion
We thus have three possible scenarios for generating long term volatility from metaorder impact alone. The idea that metaorders are correlated between one another, with roughly the same long memory as within each one of them, seems plausible, is compatible with available data and provides a natural extension of the propagator scenario: diffusive prices emerge from the subtle interplay between decaying impact and autocorrelated order flow. Such a scenario predicts a weak dependence of the prefactor $Y$ of the square-root law with the participation rate of the metaorder, as $\widetilde{\varphi}^{\beta}$ with $\beta\approx 0.2$.
@@
Finally, long term volatility may result from the random component of impact, assumed to be permanent. This bypasses the paradox that average impact decays and, in the absence of correlations between metaorders, should lead to subdiffusion. In such a scenario, volatility is induced by trading activity alone, even when average impact is zero. Still, the fact that impact fluctuations are proportional to average impact, as postulated in Eq. (68), is important to recover the correct relation between peak impact $\mathcal{I}(Q)$ and $\sigma\sqrt{Q/\phi}$, Eq. (73), which is now independent of $\widetilde{\varphi}$.
We now turn to the study of the covariance of price changes and volume imbalances, to see whether we can constrain the theory further. In fact, we will see that the data strongly favours an interpretation based on the average impact of long-range correlated metaorders. The last scenario, where volatility arises from the fluctuating part of impact, does not pass the test, as it is unable to reproduce the detailed structure of the correlations between price changes and generalized volume imbalances.
@@ 5 Covariance between order flow imbalance and prices changes
An interesting quantity that can be computed within our model and easily measured empirically using the public tape of trades and prices is the “aggregated” impact $\mathbb{E}[\Delta|I^{a}]$, conditioned to a certain value of imbalance $I^{a}$, as studied in [plerou2002quantifying, patzelt2018universal]. We know that such a quantity behaves very differently from the square-root law, and has non-trivial scaling properties as a function of $T$, see [patzelt2018universal] for $a=0$ and $a=1$. For $a=0$, in particular, one finds that the initial slope of $\mathbb{E}[\Delta|I^{0}]$ as a function of $I^{0}$ scales like $T^{-\omega}$ with $\omega\approx 1/4$ [patzelt2018universal, bouchaud2018trades], a result we confirm in Fig. 7.
@@ 5.5 With “informed” metaorders
Up to now, we assumed that there are no correlations between the sign of metaorders $\varepsilon$ and the “Fundamental” component of price changes on time scale $T$, $\sigma_{F}\xi\sqrt{T}$, see Eq. (68). If we rather assume that $\mathbb{E}[\varepsilon\xi]=\rho/\sqrt{\nu T}$, where $\rho$ measures the average amount of information of individual metaorders, [^12] For a detailed discussion of the scaling with $T$, see [bouchaud2010impact], section 16.1.3. The idea is that out of $N_{T}=\nu T$ metaorders of random signs, an excess fraction $\sim\sqrt{\nu T}$ is possibly informed. It is important to stress that, in the spirit of the Kyle model [kyle1985continuous], the fundamental component $\sigma_{F}\xi\sqrt{T}$ is not a mechanical consequence of impact. we find an extra contribution to $\mathbb{E}[\Delta_{T}\cdot I_{T}^{a}]$ computed above, which reads:
@@ 5.6 The correlation coefficient
For $a\to 0$, the power of $T$ is very close to zero for our default choice of parameters. For higher values of $a$, the exponent becomes positive and therefore this off-diagonal contribution adds an increasing function of $T$ [^13] In the asymptotic $T\to\infty$ limit, one should take into account the fact that $\Sigma_{a}$ itself becomes dominated by the off-diagonal contribution, see the discussion around Eq. (39). Hence the correlation does saturate when $T\to\infty$, as it should be.
@@ 5.7 Empirical data
    Provided the main source of price moves is the average impact of random metaorders, the covariance $\operatorname{\mathbb{E}}[\Delta_{T}\cdot I_{T}^{a}]$ behaves as a power-law of $T$ with an effective exponent that is non-monotonic in $a$, reaching a minimum for some value of $a$, see Eq. (80). If instead volatility is dominated by the random component of impact (Eq. (68)) with a “Fundamental” component $\sigma_{F}\xi$, a linear behaviour in $T$ independent of $a$ should be observed;
@@ 5.7.4 Empirical covariance: conclusion
  • Stocks and futures seem to differ quantitatively when it comes to the correlation between order imbalance and price changes. In particular, the correlation between the two is stronger for futures contract, and is reached for longer time intervals $T$ and with more weight given on large child order volumes (i.e. $a^{\star}=1$ instead of $a^{\star}=1/2$). This suggests a stronger role of metaorder correlations for futures than for stocks.
  • The hypothesis according to which most of the volatility comes from fundamental information is hard to reconcile with the data. First, as discussed in section 4.5, this assumption does not enable one to rationalize why the square-root law, which applies to all metaorders (informed or not [gomes2015market, bucci2018slow, sato2024does]) is proportional to volatility. Second, it predicts that the covariance of price returns and volume imbalances is proportional to $T$ independently of $a$, at odds with empirical data. As argued in [bouchaud2006random, taranto2018linear, bouchaud2018trades], the most plausible hypothesis is that volatility stems from trading alone. In other words, the excess volatility puzzle has a microstructural origin.